# SceneVerse: Scaling 3D Vision-Language Learning for Grounded Scene Understanding

### v1

@@ Abstract
3D vision-language grounding, which focuses on aligning language with the 3D physical environment, stands as a cornerstone in the development of embodied agents. In comparison to recent advancements in the 2D domain, grounding language in 3D scenes faces several significant challenges: (i) the inherent complexity of 3D scenes due to the diverse object configurations, their rich attributes, and intricate relationships; (ii) the scarcity of paired 3D vision-language data to support grounded learning; and (iii) the absence of a unified learning framework to distill knowledge from grounded 3D data. In this work, we aim to address these three major challenges in 3D vision-language by examining the potential of systematically upscaling 3D vision-language learning in indoor environments. We introduce the first million-scale 3D vision-language dataset, SceneVerse, encompassing about $68$K 3D indoor scenes and comprising $2.5$M vision-language pairs derived from both human annotations and our scalable scene-graph-based generation approach. We demonstrate that this scaling allows for a unified pre-training framework, Grounded Pre-training for Scenes (GPS), for 3D vision-language learning. Through extensive experiments, we showcase the effectiveness of GPS by achieving state-of-the-art performance on all existing 3D visual grounding benchmarks. The vast potential of SceneVerse and GPS is unveiled through zero-shot transfer experiments in the challenging 3D vision-language tasks.
@@
$*{}^{*}$ indicates equal contribution.
1 Introduction
The foundation of human cognitive development lies in the grounding of language within the physical world [73, 46, 97]. Recent progress in Large Language Models (LLMs[11, 75], often referred to as “foundation models” [10], has markedly promoted the alignment between vision and language [66, 3, 51] through utilizing billion-scale vision-language datasets [71, 96]. Nonetheless, with these advancements predominantly focusing on the 2D domain, the grounded understanding of 3D physical environments remains in an incipient stage [16, 1, 5]. Recognizing the pivotal role of grounded 3D experiences in shaping human cognition [7, 8] and the delayed research development in this area, there is a compelling need to intensify the exploration into the vision-language learning challenge, specifically in the context of 3D scenes.
@@ 1 Introduction
Seeking insights from the 2D vision-language (2D-VL) achievements, a major factor to the success was the notable scale-up of paired vision-language data [45, 15, 71]. However, applying these principles directly from 2D to 3D is fraught with challenges. Primarily, 3D data collection heavily relies on the scanning device, making it inherently much more complex and expensive than gathering 2D images. Despite steady efforts to increase the volume of 3D scene data [23, 87, 58, 9], most datasets remain limited to thousands of scenes, substantially lagging behind the scale of existing 2D datasets. This gap is further widened by the inherent complexities of 3D scenes, which feature a multitude of object instances with diverse attributes, varying arrangements, and intricate inter-object relationships. These unique aspects of 3D scenes not only make the accurate description of objects and their relations more challenging but also considerably increase the number of language descriptions required for thorough scene depiction. Consequently, this presents a significant challenge in obtaining a sufficient supply of high-quality paired scene-language data crucial for grounded scene understanding.
@@
To confront these challenges, we propose consolidating current efforts to build up SceneVerse, the first million-scale dataset aimed at advancing 3D vision-language (3D-VL) learning for grounded scene understanding. At the scene level, we unify 3D scene data from existing datasets [23, 78, 9, 67, 40] and supplement the collection with synthetic scenes [95, 27]. This compilation represents the most extensive 3D scene data gathered to date, amounting to $68,406$ scenes for grounding. Additionally, we propose an automated generation pipeline utilizing 3D scene graphs [4, 79] and LLMs to create comprehensive, high-quality scene-language pairs. This refined collection, including $190,836$ human annotated pairs and totaling $2.5$M scene-language pairs, provides detailed and comprehensive portrayals of both object-level and scene-level descriptions within the 3D scene.
@@
We thoroughly investigate the potential offered by the data scale-up in SceneVerse with large-scale pre-training. Specifically, we present a novel and unified pre-training framework, Grounded Pre-training for Scenes (GPS), which is designed with scene-level and object-level alignment objectives and devoid of auxiliary losses and designs. Through multi-level contrastive alignment, we observe significant performance improvements across all existing 3D visual grounding benchmarks, achieving new state-of-the-art results through a simple and effective pre-training process. Additionally, we unveil the vast possibilities offered by SceneVerse and GPS in 3D-VL tasks in a zero-shot transfer setting. At last, we provide a more comprehensive understanding of the data-scaling effects in SceneVerse through extensive ablative experiments to point out future directions.
Our main contributions can be summarized as follows:
@@
  • We propose GPS, an efficient transformer-based model trained with multi-level scene-text alignment that achieves state-of-the-art results on all existing 3D-VL grounding benchmarks, benefiting from pre-training on multi-level scene-language pairs in SceneVerse.
  • We demonstrate that with the data scale-up and model design, our pre-trained models exhibit emerging zero-shot generalization capabilities in grounded scene understanding, paralleling the successes seen in 2D-VL models.
@@ 2 Related Work
Obtaining aligned 3D-language data is an inherently difficult task. In 3D object modeling, pioneering works like ShapeNet [14] sourced 3D assets from online repositories, leading to follow-up proliferation of high-quality 3D object datasets [60, 22, 81]. Notably, recent developments include internet-scale data collection with Objaverse [26, 25], accompanied by the integration of object-level captions [83] for 3D-language alignment. Models trained on these datasets demonstrate an enhanced understanding of objects, evident in classification [52], generation [53], and captioning tasks [55].
In contrast, developing datasets for grounded 3D scene understanding is even more challenging due to the extensive requirements for scene acquisition and annotation. Existing works curate RGB-D and scanned indoor scene datasets [23, 13, 78, 9, 67, 58] initially used for benchmarking classical grounding tasks like 3D object detection and segmentation [30, 42, 59, 77, 72]. These semantically labeled scenes are subsequently applied in fine-grained scene grounding tasks like object referral [16, 1, 93], captioning [20, 88, 17, 19], vision-language-navigation [56, 80, 63, 38] and reasoning [5, 57, 37]. Recent work exploits the representation of 3D scene graphs (3DSGs) [4, 79, 69], which concisely describes scenes with hierarchical structures. This representation is notably advantageous for planning [2, 68] and captioning [33], owing to its compatibility with LLMs. Nevertheless, as shown in Tab. 1, these datasets are significantly constrained in both scene and language scales, underscoring the need for scaling up fine-grained scene-language-aligned data to enhance grounded scene understanding.
@@
Recent years have witnessed tremendous progress in 2D vision-language learning [76, 66, 70, 3, 49, 51, 24], empowered by transformer-based pre-training models [28, 11, 62] and large-scale image-language datasets [15, 71]. A central theme across language and 2D-VL domains is the effectiveness of data scaling [43], as demonstrated by improved alignment and expanded capabilities in open-vocabulary understanding [50, 47, 32, 44] through a simplified contrastive pre-training pipeline [66].
However, in grounded scene understanding, the primary challenge for models has been the limited availability of paired 3D scene-language data, which restricts the application of insights gained from 2D-VL. Current models for 3D scene grounding [35, 94, 86, 6, 54, 40, 41, 18, 82] heavily rely on task-specific knowledge in both model and loss designs or advanced optimization strategies [98]. To bridge this gap, there has been a growing emphasis on employing pre-trained 2D-VL models for 3D-VL [91, 34, 83, 36, 92, 64, 74]. Nonetheless, these models predominantly draw on information available from 2D-VL models (e.g., object attribute, affordance, etc.), falling short on capturing crucial information like object spatial relationships, which are only attainable through 3D data. This urges the need for a multi-level alignment between language and 3D scenes, particularly regarding 3D-specific information. Considering the nascent stage of existing 3D pre-training methods [29, 84, 98], we believe SceneVerse and GPS have the potential to spearhead new avenues in 3D-VL research.
Figure 2: SceneVerse collection and statistics. Given a 3D scene (a), our automated pipeline (c) generates three types of description including scene caption, object caption and object referral. (b) The comparison of different language sources and data composition.
Table 1: Comparison of SceneVerse with existing 3DVL Datasets. SceneVerse expands the data scale of prior work by an order of magnitude. Anno.: human annotations. Syn.: template or LLM generated descriptions.
Dataset 3D Data Language Total
Scene Object Anno. Syn.
ScanRefer[16] |||| |||| 52K - 52K
ReferIt3D[1] 42K 200K 242K
ScanQA[5] 1.5K 33K 27K - 27K
SQA3D[57] |||| |||| - 33K 33K
Multi3DRefer[93] 52K 10K 62K
Cap3D[55] -  [^6] K 58K  [^6] K 724K
ScanScribe[98] 3K 56K 94K 184K 278K
SceneVerse 68K 1.5M 190K 2.3M 2.5M
@@ 3 SceneVerse
SceneVerse is the first million-scale dataset designed for grounded scene understanding. Our 3D scenes are curated from diverse existing datasets of both real and synthetic environments. Harnessing the power of 3D scene graphs and LLMs, we introduce an automated pipeline to generate comprehensive and high-quality language for both object-level and scene-level descriptions. We additionally incorporate the most extensive human-annotated object referrals to date, providing new training sources and benchmarks in this field.
@@ 3.1 Scene Curation
To address the scarcity of available 3D scene data, we construct SceneVerse by unifying 3D scene data from various existing datasets. We use real-world scene datasets, including ScanNet [23], ARKitScenes [9], HM3D [67], 3RScan [78] and MultiScan [58], alongside synthetic environments from Structured3D [95] and ProcTHOR [27]. The inclusion of these synthetic datasets is mainly motivated by their potential as scalable data sources for 3D-VL alignment. To ensure cohesion across various sources, we conduct preprocessing steps such as room segmentation, point subsampling, axis alignment, normalization, and semantic label alignment. Each scan is represented by a point cloud $\mathrm{P}\in\mathbb{R}^{N\times 8}$, wherein each point is defined by its 3D coordinates, RGB color, instance id and semantic label. In total, we curate $68,406$ 3D scenes in SceneVerse.
@@
Our 3D scene graph is defined as a set of tuples $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the nodes $\mathcal{V}$ comprises $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level. Each node $v$ represents one distinct 3D object instance, parameterized by its centroid ${\bm{p}}_{i}\in\mathbb{R}^{3}$ and bounding box size of ${\bm{b}}_{i}=(b_{x},b_{y},b_{z})\in\mathbb{R}^{3}$. The edges $\mathcal{E}$ represent spatial relationships between nodes.
To construct the scene graph $\mathcal{G}$, we first instantiate the nodes with the instance annotation from the point clouds and assign object classes with their corresponding semantic labels. Following prior work[1, 79], we consider the following spatial relations.
Vertical proximity
This encompasses both in-contact relationships (e.g., support, inside, embed), and non-contact ones (e.g., above, below).
Horizontal proximity
Horizontal relationships describe the proximity relations like in front of, next to, behind, etc. Relationships like left, right are contextually dependent on a reference view, where another anchor object is utilized to establish the view direction. The distance between the two objects is also calculated to describe whether the objects are far or near in space.
Multi-object Relationships
This models the spatial arrangement of multiple objects, e.g., align and between.
The node hierarchy is decided by the support relationship. We traverse all the object nodes to calculate spatial relationships, which undergo an automatic verification procedure to rectify incorrect ones. For a more detailed description of the scene graph construction and relationship determination, please refer to Sec. A.2.
@@ 3.3 Language Generation with LLMs
The scene-language pairs in SceneVerse aim to capture varying aspects of the 3D scene, which include detailed object attribute descriptions in object captioning, spatial relationships between objects in object referral, and global scene descriptions in scene captioning. Based on the 3D scene graph, we utilize both templates and LLMs to automatically generate descriptions on these three granularities.
@@
Object captions aim to provide detailed descriptions of an object’s visual and physical properties, facilitating object-level grounding with its distinctive features. Given the multi-view images, we utilize the point cloud of the object $v\in\mathcal{V}$ to identify its occurrence in the images through rendering. The images are then cropped with the rendered bounding boxes and processed through BLIP2 [48] to generate initial object captions. To refine the captions, we select the top 10 sentences with the highest CLIP [66] similarity score and minimal occlusion. The selected sentences are fed into a LLM to obtain a coherent summary of the object captions. In this process, we explicitly instruct the language model to identify and correct the potential errors. The detailed object captioning pipeline is illustrated in Sec. A.3.
@@
The scene-level captions emphasize global information, portraying the key objects in the scene along with their attributes and functionalities. We leverage the constructed 3D scene graph and prompt LLMs to generate these captions. To enhance the diversity of scene captions, we utilize a subgraph sampling strategy, where a subset of edges and nodes are randomly sampled as the scene context. The object counts are also provided as LLM prompts, together with the room type and object attributes if such annotations are available in the dataset.
3.4 Referral Annotation by Humans
In addition to automatically generated scene-text pairs, SceneVerse includes the most comprehensive set of human-annotated, context-rich object referrals to date, serving as a valuable benchmark for assessing grounded scene understanding capabilities. The human annotations contain $96,863$ descriptions in ARKitScenes [9], HM3D [67] and MultiScan [58]. During the annotation process, one human annotator was assigned to write at least 20 words to distinctly refer to a single 3D object within a 3D scene. Each referral text then undergoes independent verification by two additional reviewers, both mandated to accurately locate the referenced object based on the 3D scene and the annotated referral text. Any object referrals that do not pass the verification by either reviewer are flagged for re-annotation.
3.5 Details and Statistics
In total, SceneVerse comprises a total of $68,406$ room-level 3D scans, with the source composition shown in Fig. 2 (b). The dataset contains $1.5$M object instances, comprising $21$ types of relationships following prior work [1, 79]. For the language descriptions, we generate $1$M template-based texts and $1$M sentences by LLM rephrased by Llama [75] and GPT-3.5 [61]. All the rephrasing and summary prompts, along with the complete set of relationships, are detailed in Sec. A.3. To verify the efficacy of our automated language generation pipeline, we conduct a quality check (QC) where 12K generated object-level descriptions are randomly selected for human verification, achieving a $96.93\%$ pass rate. This shows the capability of our proposed scene-graph-based generation approach to produce high-quality language descriptions, laying a robust foundation for future scalability.

### v3

# State Key Laboratory of General Artificial Intelligence, BIGAI %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.https://scene-verse.github.io

@@ Abstract
3D vision-language (3D-VL) grounding, which aims to align language with 3D physical environments, stands as a cornerstone in developing embodied agents. In comparison to recent advancements in the 2D domain, grounding language in 3D scenes faces two significant challenges: (i) the scarcity of paired 3D-VL data to support grounded learning of 3D scenes, especially considering complexities within diverse object configurations, rich attributes, and intricate relationships; and (ii) the absence of a unified learning framework to distill knowledge from grounded 3D data. In this work, we aim to address these major challenges in 3D-VL by examining the potential of systematically upscaling 3D-VL learning in indoor scenes. We introduce the first million-scale 3D-VL dataset, SceneVerse, encompassing $68$K indoor scenes and comprising $2.5$M vision-language pairs collected from both human annotations and our scalable scene-graph-based generation approach. We demonstrate that this scaling allows for a unified pre-training framework, Grounded Pre-training for Scenes (GPS), for 3D-VL learning. Through extensive experiments, we showcase the effectiveness of GPS by achieving state-of-the-art performance on existing 3D visual grounding and question-answering benchmarks. We also show that the data scaling effect is not limited to GPS, but is generally beneficial for models on tasks like 3D semantic segmentation. The vast potential of SceneVerse and GPS is unveiled through zero-shot transfer experiments in challenging 3D-VL tasks.
Keywords:
3D Vision-Language Data Scaling Grounded Scene Understanding
1 Introduction
The foundation of human cognitive development lies in the grounding of language within the physical world [84, 55, 111]. Recent progress in Large Language Models (LLMs[11, 86, 10] has markedly promoted the alignment between vision and language [78, 3, 61] utilizing billion-scale vision-language datasets [82, 110]. However, with these advancements predominantly focusing on the 2D domain, the grounded understanding of 3D physical environments remains in an incipient stage [16, 1, 5]. Recognizing the pivotal role of grounded 3D experiences in shaping human cognition [7, 8], there is a compelling need to focus on exploring vision-language learning in the context of 3D scenes.
@@ 1 Introduction
Seeking insights from success in 2D vision-language (2D-VL) learning, a major factor to the success was the notable scale-up of paired vision-language data [54, 15, 82]. However, applying this experience directly from 2D to 3D is fraught with challenges. Primarily, 3D data collection heavily relies on the scanning device, making it inherently much more complex and expensive than gathering 2D images. Despite steady efforts to increase the volume of 3D scene data [25, 101, 68, 9], most datasets remain limited to thousands of scenes, substantially lagging behind the scale of existing 2D datasets. This gap is further widened by the inherent complexities of 3D scenes, which feature a multitude of object instances with diverse attributes, varying arrangements, and intricate inter-object relationships. These unique aspects of 3D scenes not only make the accurate description of objects and their relations more challenging but also considerably increase the number of language descriptions required for thorough scene depiction. Consequently, this presents a significant challenge in gathering sufficient and high-quality paired scene-language data for grounded scene understanding.
@@
To confront these challenges, we propose SceneVerse, the first million-scale dataset aimed at advancing 3D vision-language (3D-VL) learning for grounded scene understanding. At the scene level, we unify 3D scene data from existing datasets [25, 88, 9, 79, 68], aligning scenes and annotations from various capturing sources, and supplement the collection with synthetic scenes [109, 29]. This compilation represents the most extensive 3D scene data gathered to date, amounting to $68$K scenes. For language, we first present $97$K newly-annotated referring expressions, the most extensive thus far. We additionally propose an automated generation pipeline utilizing 3D scene graphs [4, 89] and LLMs to create comprehensive, high-quality scene-language pairs. This refined collection, totaling $2.5$M scene-language pairs, provides detailed and comprehensive descriptions of both object-level and scene-level descriptions within the 3D scene.
@@
We thoroughly investigate the potential offered by SceneVerse with large-scale pre-training, introducing Grounded Pre-training for Scenes (GPS), a novel and unified pre-training framework designed for scene and object-level alignment without auxiliary losses. Through multi-level contrastive alignment, we achieve significant performance boosts on 3D-VL tasks, such as grounding and question answering, setting new state-of-the-art results via a simple and effective pre-training process. We unveil the vast possibilities offered by SceneVerse and GPS in 3D-VL tasks in a zero-shot transfer setting. Additionally, we show that the scaling effect in SceneVerse is not limited to GPS, but generally benefits models in tasks like 3D semantic segmentation. At last, we offer deeper insights into the data-scaling in SceneVerse through extensive ablative experiments, pointing out future directions. Our main contributions are as follows:
  1. 1.
@@
  2. 2.
    We propose GPS, a transformer-based model trained with multi-level scene-text alignment that achieves state-of-the-art results on existing 3D-VL grounding and question-answering benchmarks by pre-training on SceneVerse.
  3. 3.
    We demonstrate that with the data scale-up and model design, our pre-trained GPS exhibit emerging zero-shot generalization capabilities in grounded scene understanding. We also show that this scaling effect is not limited to GPS, but is generally beneficial for models on tasks like semantic segmentation.
@@ 2 Related Work
Obtaining aligned 3D-language data is a inherently difficult task. In 3D object modeling, pioneering works like ShapeNet [14] sourced 3D assets from online repositories, leading to a proliferation of high-quality 3D object datasets [71, 24, 94]. Notably, recent developments include internet-scale data collection with Objaverse [28, 27], accompanied by the integration of object-level captions [96] for 3D-language alignment. Models trained on these datasets demonstrate an enhanced understanding of objects, evident in classification [62], generation [63], and captioning tasks [65].
In contrast, developing datasets for grounded 3D scene understanding is even more challenging due to the extensive requirements for scene acquisition and annotation. Existing works curate RGB-D and scanned indoor scene datasets [25, 13, 88, 9, 79, 68] and synthetic scenes [29, 109, 52, 98] used for benchmarking tasks like 3D object detection and segmentation [32, 48, 69, 87, 83]. These semantically labeled scenes are subsequently used in fine-grained scene grounding tasks like object referral [16, 1, 107], captioning [22, 102, 17, 19], vision-language-navigation [66, 91, 75, 41] and reasoning [5, 67, 42]. Recent works exploit the representation of 3D scene graph (3DSG) [20, 4, 89, 81], which concisely describes scenes with hierarchical structures. This representation is notably advantageous for planning [2, 80] and captioning [37], owing to its compatibility with LLMs for flexible description generation [42, 43]. Nonetheless, as shown in Tab. 1, most datasets are constrained in both scene and language scales, underscoring the need for scaling up fine-grained and aligned scene-language data to enhance grounded scene understanding.
@@
Recent years have witnessed tremendous progress in 2D-VL [78, 3, 59, 61, 26, 21], empowered by transformer-based pre-training models [30, 11, 74] and large-scale image-language datasets [15, 82]. A central theme across 2D-VL domains is the effectiveness of data scaling [51], as demonstrated by improved alignment and expanded capabilities in open-vocabulary understanding [60, 56, 34, 53] through a contrastive pre-training pipeline [78].
However, in grounded scene understanding, the primary challenge for models has been the limited availability of paired 3D scene-language data, which restricts the application of insights drawn from 2D-VL. Current models for 3D scene grounding [39, 108, 100, 6, 64, 45, 47, 18, 95] heavily rely on task-specific knowledge in both model and loss designs or advanced optimization strategies [112]. To bridge this gap, there has been a growing emphasis on employing pre-trained 2D-VL models for 3D-VL [105, 38, 96, 40, 106, 76, 85]. Yet, these models mostly draw information available from 2D-VL models (e.g., object attribute, affordance, etc.), falling short on capturing crucial 3D information like object spatial relationships which are necessary for more fine-grained tasks such as grounded human-scene [50, 92, 46, 49, 93, 20] and robot-scene interactions modeling [72, 57, 70, 35]. This urges the need for a multi-level alignment between language and 3D scenes, particularly regarding 3D-specific information. Considering the nascent stage of existing 3D pre-training methods [31, 97, 112, 113], we believe SceneVerse and GPS have the potential to spearhead new avenues in 3D-VL research.
Table 1: Comparison of SceneVerse with existing 3DVL Datasets. SceneVerse expands the data scale of prior work by order of magnitude. “VG” stands for Visual Grounding, “QA” for Question Answering, “PT” for Pre-training and “MT” for Multi-tasking. “Anno.” denotes language from human annotations and “Syn.” for template-based or LLM generated descriptions.
Dataset 3D Task Obj. Scene Obj. Quality New Existing Total
Scene Obj. Caption Caption Referral Check Anno. Syn. Anno. Syn.
ScanRefer [16] |||| |||| VG 52K - - - 52K
ReferIt3D [1] VG 42K 200K - - 242K
ScanQA [5] 1.5K 33K QA - - - 27K - - - 27K
SQA3D [67] |||| |||| QA - - - 33K - - - 33K
Multi3DRefer [107] VG - 10K 52K - 62K
Cap3D[65] -  [^6] K VG 58K  [^6] K - - 724K
ScanScribe [112] 3K 56K PT - 90K 94K 94K 278K
3D-LLM [42] 1.5K 186K MT - 659K - - 659K
EmbodiedScan [90] 5K 890K VG - 970K - - 970K
LEO [43] 3K 56K MT - 188K 235K 90K 513K
SceneVerse 68K 1.5M VG 96K 2.1M 94K 200K 2.5M
@@ 3 SceneVerse
SceneVerse is designed for grounded scene understanding with 3D scenes curated from diverse existing datasets of both real and synthetic environments. Regarding language, we employ both human annotation and a novel automated generation pipeline to collect comprehensive and high-quality language for both object-level and scene-level descriptions. We provide details regarding data collection in the following sections.
@@ 3.1 Scene Curation
To address the scarcity of available 3D scene data, we construct SceneVerse by unifying 3D scene data from various existing datasets. We use real-world scene datasets, including ScanNet [25], ARKitScenes [9], HM3D [79], 3RScan [88] and MultiScan [68], alongside synthetic environments from Structured3D [109] and ProcTHOR [29]. The inclusion of these synthetic datasets is mainly motivated by their potential as scalable data sources for 3D-VL alignment. To ensure cohesion across various sources, we conduct preprocessing steps such as room segmentation, point subsampling, axis alignment, normalization, and semantic label alignment. Each scan is represented by a point cloud $\mathrm{P}\in\mathbb{R}^{N\times 8}$, wherein each point is defined by its 3D coordinates, RGB color, instance id, and semantic label. In total, we curate $68,406$ 3D scenes in SceneVerse.
3.2 Referral Annotation by Humans
In the curated scenes of SceneVerse, we present the most comprehensive set of human-annotated, context-rich object referrals to date, serving as a valuable benchmark for assessing grounded scene understanding capabilities. The human annotations contain $96,863$ descriptions in ARKitScenes [9], HM3D [79] and MultiScan [68]. During the annotation process, one human annotator was assigned to write at least 20 words to distinctly refer to a single 3D object within a 3D scene. Each referral text then undergoes independent verification by two additional reviewers, both mandated to accurately locate the referenced object based on the 3D scene and the annotated referral text. Any object referrals that do not pass the verification by either reviewer are flagged for re-annotation.
Figure 2: SceneVerse collection and statistics. Given a 3D scene (a), our automated pipeline (c) generates three types of description including scene caption, object caption and object referral. (b) SceneVerse data comparison and composition.
@@ 3.2 Referral Annotation by Humans
Our 3D scene graph is defined as a hierarchical graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Each node $v\in\mathcal{V}$ represents one distinct 3D object instance, parameterized by its centroid ${\bm{p}}_{i}\in\mathbb{R}^{3}$ and bounding box size of ${\bm{b}}_{i}=(b_{x},b_{y},b_{z})\in\mathbb{R}^{3}$. The edges $\mathcal{E}$ represent spatial relationships between nodes. To construct the scene graph $\mathcal{G}$, we instantiate the nodes with the instance annotation from the point clouds and assign object classes with their corresponding semantic labels. Following prior work[1, 89], we consider the Vertical proximity, Horizontal proximity and Multi-object Relationships as spatial relations. For a more detailed description of the scene graph construction and relationship determination, please refer to supplementary.
@@ 3.4 Language Generation with LLMs
The scene-language pairs in SceneVerse aim to capture varying aspects of the 3D scene, including detailed object attributes in object captioning, spatial relationships between objects in object referral, and global scene descriptions in scene captioning. Based on 3D scene graphs, we utilize both templates and LLMs to automatically generate descriptions on these three granularities.
@@
Object captions aim to provide detailed descriptions of an object’s visual and physical properties, facilitating object-level grounding with its distinctive features. Given the multi-view images, we utilize the point cloud of the object $v\in\mathcal{V}$ to identify its occurrence in the images through rendering. The images are then cropped with the rendered bounding boxes and processed through BLIP2 [58] to generate initial object captions. We select the top 10 sentences with the highest CLIP [78] similarity score and minimal occlusion and utilize an LLM to obtain a refined and coherent summary of the captions. The detailed object captioning pipeline is illustrated in supplementary.
@@
The scene-level captions emphasize global information, portraying the key objects in the scene along with their attributes and functionalities. We use the constructed 3D scene graph and prompt LLMs to generate these captions. We random sample a subset of edges and nodes from the scene graph each time as the scene context to enhance the diversity of scene captions. The object counts are also provided as LLM prompts, together with the room type and object attributes if such annotations are available in the dataset.
3.5 Data Quality and Statistics
Data Quality
The $96$K human-annotated set of SceneVerse is collected through AMT, where $82$ humans are employed for annotation and $18$ for verification. All final annotations passed the reference verification, with a re-annotation rate of $4.8\%$. For our automatic language generation pipeline, we conduct extensive prompt tuning and iterate with human feedback for LLMs on object captioning, summary, and rephrasing. To verify the efficacy of the pipeline, we conduct a quality check where $12$K generated object-level descriptions are randomly selected for human verification. Results demonstrate a $96.93\%$ pass rate, surpassing that in ReferIt3D [1] with $86.1\%$ pass rate on $2$K samples.
Statistics
In total, SceneVerse comprises a total of $68,406$ room-level 3D scans, with the source composition shown in Fig. 2 (b). The dataset contains $1.5$M object instances ranging in 2290 object categories. Our generated 3D scene graph comprises $21$ types of relationships following prior work [1, 89]. For the language descriptions, we generate $1$M template-based texts and $1$M sentences rephrased by Llama [86] and GPT-3.5 [73]. As can be seen from the radar chart and examples in Fig. 2, the diversity of the LLM-refined data, particularly in sentence length and variety, closely aligns with the characteristics of annotated descriptions, surpassing the template-based data. Together with existing sources (294K) and our newly annotated set (96K), SceneVerse contains $2.5$M scene-language pairs in total. All the rephrasing and summary prompts, along with the complete set of relationships, are detailed in supplementary.